OpenStudy (anonymous):

Find the volume of the region inside z^2=x^2+y^2 between z=1 and z=9. Is the z limits 1 < z < 9... and not really sure on the others. Seems simple but I'm confused.

OpenStudy (amistre64):

you have a sphere with a section between 1 and 9 is what it looks like to me

OpenStudy (amistre64):

its a cone aint it

OpenStudy (anonymous):

so 1<z<9, 1<y<3 and -rad(z^2-y^2)<x<rad(z^2-y^2) I assume that is wrong...

OpenStudy (amistre64):

dunno yet, we will determine that after I get a good bead on it

OpenStudy (amistre64):

spose x=0; the z^2 = y^2 z = +- sqrt(y^2) = +- y

OpenStudy (anonymous):

and it is a cone

OpenStudy (amistre64):

same calcs for the y=0 parts and we get a cone about the z axis

OpenStudy (amistre64):



OpenStudy (anonymous):

is it easiest to use polar coordinates or cartesian?

OpenStudy (amistre64):

for this, either or ... since the radius of each circle is defined by x or y; we can simply say: \[A=\pi [f(x)]^2\] or f(y) ^2 but we have to pick one of them

OpenStudy (amistre64):

cylindrical coords is an option too; but they all amount to the same thing i believe

OpenStudy (amistre64):

the radius moves from x=0 to x=sqrt(z)

OpenStudy (amistre64):

\[\int_{1}^{9}\int_{0}^{\sqrt{z}}\ 2\pi\ x dx.dz\] should do it

OpenStudy (anonymous):

I understand the limits. But why do you have 2*pi*x? Sorry I'm new to this.

OpenStudy (amistre64):

quite frankly, i had to adjust my original thoughts :) essentially this adds up all the circumferences from the center out; the sum of all the circumferences add up to the area; then we do for all the points along z from 1 to 9

OpenStudy (amistre64):

 add up all the circumferences to get area of 1 circle

OpenStudy (amistre64):

the area of each circle is: \(z\ \pi \) int z = z^2/2 from 1 to 9 = 9^2/2 - 1^2/2

OpenStudy (amistre64):

so 40?

OpenStudy (amistre64):

40 pi :)

OpenStudy (anonymous):

I get the way you figured it out but how would you formally set it up as a double integral.

OpenStudy (amistre64):

one way is the way I did it; im sure there are other ways of course

OpenStudy (amistre64):

we could define a triple as the radius, the angle, and the z axis

OpenStudy (amistre64):

\[\int_{1}^{9}\int_{0}^{2pi}\int_{0}^{\sqrt{z}} dr.d\theta .dz\]

OpenStudy (anonymous):

That makes perfect sense. Thanks!! It should be "r*dr*d(theta)*d(z)" though right?

OpenStudy (amistre64):

i believe it starts with the function r=1

OpenStudy (anonymous):

Could you explain that? I don't understand what you mean it stats with the function r=1.

OpenStudy (amistre64):

lets try it : dr = r, sqrt(z) - 0 = sqrt(z) : sqrt(z) dt = sqrt(z) t ; 2pi sqrt(z) - 0 : 2pi sqrt(z) dz = 4pi/3 z^(3/2) hmmm, i could be wrong :) but yeah, it works out something along these lines integration is just adding up alot of points, so as long as you know what your adding up and how it moving; youre good to go

OpenStudy (amistre64):

the circumference of a circle = 2pi r , say we want a radius of 5 \[\int_{0}^{2pi}\int_{0}^{5}dr.dt=\int_{0}^{2pi}r;[0,5]\ dt\to\ \int_{0}^{2pi}5dt\] \[\int_{0}^{2pi}5dt\to\ 5t;[0,2pi] = 10pi\] ok, so that route gets us to circumferece in the end

OpenStudy (amistre64):

oh well; id stick with a double of 2pi r; r=0 to sqrt(z) and z goes from 1 to 9

OpenStudy (anonymous):

Thank you, sir.

OpenStudy (amistre64):

yw